# Investigating and Defending Shortcut Learning in Personalized Diffusion Models

### v3

@@ 4.1 Adversarial Perturbation Causes Latent-space Image-Prompt Mismatch
$\mathcal{L}_{db}^{adv}(\mathbf{x}_{0},\bm{c}^{ID},\bar{\bm{x}}_{0},\bm{c};\bm{\theta})=\mathcal{L}_{\mathrm{denoise}}\left(\mathbf{x}_{0}+\bm{\delta},\bm{c}^{ID}\right)+\lambda\mathcal{L}_{\mathrm{denoise}}\left(\bar{\bm{x}}_{0},\bm{c}\right).$ (5)
Figure 2: Zero-shot classification with CLIP-based classifier.
Based on the adversarial loss in Eq. 5, with annotation of $\bm{c}^{ID}=\bm{c}\oplus ID$ where $ID$ denotes the embedding of the unique identifier, we build the causal dependency graph in Fig 3 to represent the variable relationship during the learning of the personalized diffusion model to capture the concept of instance. From this graph, we found that there are two false correlations (colored in red) derived from the prior prompt condition $\bar{\bm{c}}$ and instance condition $ID$ to the noise $\bm{\delta}$. These two paths hinder the personalized diffusion from capturing the true relationship between the instance condition $\bm{c}^{ID}$ with its semantic identity behind $\mathbf{x}_{0}$ and can cause the model to overfit to simpler relationship $ID\rightarrow\bm{\delta}$. We empirically demonstrate this point in Figure 1 via conducting sampling from models trained on perturbed images using [4] with $ID$ as prompting condition, where we found the model falsely links the $ID$ with the noisy pattern. Furthermore, using CLIP [37] as a zero-shot classifier with labels of “noise” and “person” (shown in Figure 2), we identify that the perturbed images largely diverge from their original semantic textual concept and also its paired textual prompt $\bm{c}^{ID}$ in terms of CLIP embedding space. We hypothesize that learning on such perturbed images will create contradiction and force the models to dump that chaotic perturbation pattern into the rarely-appeared identifier token $ID$ instead of learning clean identity behind $\bm{x}_{0}$.
Figure 3: Variables dependency graph when learning personalized diffusion models.
4.2 Systematic Defending with Input Purification, Decouple Learning, and Guided Sampling
Motivated by our empirical observation, we propose two directions to enhance the robustness when learning personalized diffusion on a potentially corrupted dataset: i) weakening those spurious paths toward noisy prediction; ii) enhancing those toward prior data $\bar{\mathbf{x}}_{0}$ and true instance data $\mathbf{x}_{0}$. For the first direction, we leverage the off-the-shelf super-resolution and image restoration models to improve the data quality and remove the noise $\bm{\delta}$. For the second part, we introduce a novel noise-aware contrastive learning with negative tokens to make the model focus on those concept-related features and avoid being trapped in the local minimum of learning perturbation. Furthermore, during the sampling, we guide to generate high-quality images with simple negative prompting. We present our overall framework in Figure 4 and Alg. 1 and describe each of the proposed modules in detail as follows.
Figure 4: The framework of our method. Our method includes three steps, including input denoising, decoupling training with negative tokens, and sampling with negative prompting. Firstly, we leverage off-the-shelf image restoration models for purification. Secondly, we link the instance prompt and prior prompt with negative tokens in a contrastive way to guide the model in learning the correct mapping. Lastly, we insert a noise-free negative prompt to further boost the quality.
@@ 4.2 Systematic Defending with Input Purification, Decouple Learning, and Guided Sampling
Image Purification as Image Restoration. The first intuitive and effective approach is to directly conduct denoising on the input image and seek to remove the adversarial perturbation. To achieve that, we view the perturbed images as degraded images in the image restoration domain and first leverage off-the-shelf image restoration models to convert low-quality, noisy images into high-quality, purified ones. Specifically, we use a face-oriented model named CodeFormer [38], which is trained on facial data to restore the images based on the latent code discretization. Besides, we leverage an additional diffusion-based super-resolution model to further enhance the purification on non-face region. Compared to the previous SoTA optimization-based purification method [12] and diffusion-based purification methods [13], this simple yet effective strategy yields faithful purified images with better efficiency since it only requires a single inference.
Contrastive Decouple Learning with Noise Tokens. Furthermore, in the prompt design, we add an additional tag related to the noise pattern, such as with XX noisy pattern” aiming to decouple the learning of the concept and the unwanted noise structure. We also leverage the prior dataset that is prone to have a cleaner quality to help the model differentiate the noise pattern by adding a suffix, without XX noisy pattern” into their textual prompt. During the sampling process, we also added this suffix in prompt input to guide the model in dropping those unwanted noisy patterns.
Noise-free Sampling Process. By using classifier-free guidance (CFG) [39], with a negative prompt $c_{\text{neg}}=$noisy, abstract, pattern, low quality’, we further seek to guide the model generates high-quality images related to the learned concept. Specifically, given timestamp $t$, we perform sampling using the linear combination of the good-quality and bad-quality conditional noise estimates with $w^{\text{neg}}=7.5$:
@@
$\tilde{\bm{\epsilon}}_{\theta}\left(\mathbf{z}_{t},\mathbf{c}\right)=(1+w^{\text{neg}})\bm{\epsilon}_{\theta}\left(\mathbf{z}_{t},\mathbf{c}^{\text{ID}}\right)-w^{\text{neg}}\bm{\epsilon}_{\theta}\left(\mathbf{z}_{t},\tau_{\theta}(c_{\text{neg}})\right)$ (6)
@@ 5 Experiments
In this section, we first perform evaluations on the effectiveness, efficiency and faithfulness of the proposed purification techniques. Then, we investigate the resilience of our framework under adaptive perturbation against the purification models. Lastly, we perform systematic ablation studies and analysis to provide an in-depth exploration of each module in our framework.
@@ 5.1 Experimental Setup
Datasets and Metrics. Our experiments are mainly performed on the VGGFace2 [40] face dataset following [4, 8]. Four identities are selected from each dataset, and we randomly pick eight images from each individual and split those images into two subsets for image protection and reference. Moreover, we also visually demonstrate the purification ability of our approach on the artwork painting dataset, WikiArt [41], and the CelebA [42]. For the metric, we evaluate the generated images in terms of their semantic-related quality and graphical aesthetic quality. For the semantic-related score, we compute the cosine similarity between the embedding of generated images and reference images, which we term the Identity Matching Similarity (IMS) score. We reported the weighted averaged IMS score by employing two face embedding extractors, including antelopev2 model from InsightFace library [43] following IP-adapter [44] and VGG-Net [45] from Deepface library [46] following [4]. The IMS score is computed via a weighted sum: IMS$=\lambda\text{IMS}_{\text{IP}}+(1-\lambda)\text{IMS}_{\text{VGG}}$, where $\lambda$ is set as 0.7. For the graphical quality $Q$, we report the average of two metrics: i) LIQE [47], a SoTA general image quality assessment model (with re-normalization to $[-1,+1]$); ii) CLIP-IQAC following [8], which is based on CLIP-IQA [48] with class label. See App. A.2 for details.
@@
Purification Baselines and Perturbation Methods. For purification baselines, we consider both model-free and diffusion-based approaches. The model-free methods include ❶ Gaussian Filtering, which reduces noise and detail using a Gaussian kernel; ❷ Total Variation Minimization (TVM), which reconstructs images by minimizing the difference between original and reconstructed images while enforcing smoothness; and ❸ JPEG Compression, which reduces image file size by transforming images into a compressed format. The diffusion-based methods include ❹ (Pixel)DiffPure [14], which leverages pre-trained pixel-space diffusion models to smooth adversarial noise with small-step SDEdit process [31]; ❺ LatentDiffPure, which is developed in the paper similar as DiffPure but with LDM as a purifier (two variants w/ and w/o prompting); ❻ DDSPure [49], which finds an optimal timestamp for adversarial purification with SDEdit process; ❼ GrIDPure [15], which further conducts iterative DiffPure with small steps with grid-based splitting to improve structure similarity; and ❽ IMPRESS [12], which purifies by optimizing latent consistency with visual similarity constraints. For protective perturbation, we consider six of existing SoTA approaches, including perturbation crafted with bi-level optimization, such as FSMG, ASPL, EASPL [4], MetaCloak [8], and perturbations crafted with adversarial perturbation with fixed models, such as AdvDM [7], PhotoGuard [5], and Glaze [6]. For each setting, we set the perturbation to be ASPL by default. We set the $\ell_{\infty}$ radius to 11/255 with a six-step PGD step size of 1/255 by default following [4]. See App. A for more details.
@@
IMS $\uparrow$ Q $\uparrow$ IMS Q IMS Q IMS Q IMS Q IMS Q IMS Q
Clean 0.27 $\pm$ 0.10 0.21 $\pm$ 0.11 0.27 $\pm$ 0.10 0.21 $\pm$ 0.10 0.27 $\pm$ 0.10 0.21 $\pm$ 0.10 0.27 $\pm$ 0.10 0.21 $\pm$ 0.12 0.27 $\pm$ 0.10 0.21 $\pm$ 0.11 0.27 $\pm$ 0.10 0.21 $\pm$ 0.11 0.27 $\pm$ 0.10 0.21 $\pm$ 0.11
Perturbed -0.02 $\pm$ 0.09 -0.61 $\pm$ 0.07 -0.04 $\pm$ 0.08 -0.62 $\pm$ 0.12 -0.04 $\pm$ 0.08 -0.60 $\pm$ 0.12 0.10 $\pm$ 0.10 -0.60 $\pm$ 0.08 0.01 $\pm$ 0.09 -0.29 $\pm$ 0.09 0.17 $\pm$ 0.09 -0.13 $\pm$ 0.08 0.25 $\pm$ 0.10 -0.19 $\pm$ 0.10
Gaussian F. 0.14 $\pm$ 0.10 -0.32 $\pm$ 0.08 0.15 $\pm$ 0.10 -0.34 $\pm$ 0.06 0.14 $\pm$ 0.10 -0.33 $\pm$ 0.07 0.15 $\pm$ 0.09 -0.44 $\pm$ 0.04 0.15 $\pm$ 0.09 -0.35 $\pm$ 0.07 0.14 $\pm$ 0.10 -0.28 $\pm$ 0.08 0.12 $\pm$ 0.09 -0.24 $\pm$ 0.07
JPEG 0.11 $\pm$ 0.10 -0.46 $\pm$ 0.10 0.14 $\pm$ 0.11 -0.43 $\pm$ 0.07 0.13 $\pm$ 0.11 -0.43 $\pm$ 0.07 0.06 $\pm$ 0.11 -0.47 $\pm$ 0.12 0.01 $\pm$ 0.11 -0.01 $\pm$ 0.11 0.12 $\pm$ 0.11 0.15 $\pm$ 0.11 0.16 $\pm$ 0.10 0.24 $\pm$ 0.14
TVM 0.03 $\pm$ 0.08 -0.45 $\pm$ 0.06 0.05 $\pm$ 0.09 -0.47 $\pm$ 0.07 0.09 $\pm$ 0.09 -0.47 $\pm$ 0.06 0.06 $\pm$ 0.09 -0.50 $\pm$ 0.04 0.07 $\pm$ 0.09 -0.47 $\pm$ 0.06 0.04 $\pm$ 0.09 -0.50 $\pm$ 0.06 -0.09 $\pm$ 0.08 -0.48 $\pm$ 0.06
PixelDiffPure -0.04 $\pm$ 0.08 -0.37 $\pm$ 0.05 0.01 $\pm$ 0.08 -0.42 $\pm$ 0.06 -0.05 $\pm$ 0.09 -0.49 $\pm$ 0.05 -0.04 $\pm$ 0.08 -0.44 $\pm$ 0.06 0.00 $\pm$ 0.09 -0.42 $\pm$ 0.06 0.04 $\pm$ 0.09 -0.44 $\pm$ 0.04 -0.09 $\pm$ 0.08 -0.57 $\pm$ 0.05
L.DiffPure-$\emptyset$ 0.06 $\pm$ 0.09 -0.45 $\pm$ 0.08 -0.03 $\pm$ 0.09 -0.50 $\pm$ 0.06 0.07 $\pm$ 0.10 -0.42 $\pm$ 0.08 -0.07 $\pm$ 0.08 -0.28 $\pm$ 0.08 -0.13 $\pm$ 0.08 -0.18 $\pm$ 0.11 0.04 $\pm$ 0.08 -0.28 $\pm$ 0.08 -0.09 $\pm$ 0.08 -0.21 $\pm$ 0.08
L.DiffPure 0.02 $\pm$ 0.08 -0.36 $\pm$ 0.09 -0.02 $\pm$ 0.09 -0.52 $\pm$ 0.06 0.02 $\pm$ 0.09 -0.41 $\pm$ 0.09 -0.05 $\pm$ 0.08 -0.26 $\pm$ 0.06 -0.09 $\pm$ 0.08 -0.18 $\pm$ 0.10 -0.02 $\pm$ 0.08 -0.26 $\pm$ 0.06 -0.13 $\pm$ 0.08 -0.22 $\pm$ 0.08
DDSPure 0.17 $\pm$ 0.11 -0.23 $\pm$ 0.07 0.20 $\pm$ 0.10 -0.28 $\pm$ 0.06 0.23 $\pm$ 0.09 -0.26 $\pm$ 0.06 0.16 $\pm$ 0.10 -0.48 $\pm$ 0.07 0.13 $\pm$ 0.10 -0.30 $\pm$ 0.07 0.17 $\pm$ 0.10 -0.25 $\pm$ 0.06 0.13 $\pm$ 0.10 -0.25 $\pm$ 0.07
GrIDPure 0.17 $\pm$ 0.10 -0.11 $\pm$ 0.07 0.18 $\pm$ 0.10 -0.09 $\pm$ 0.07 0.20 $\pm$ 0.10 -0.16 $\pm$ 0.07 0.12 $\pm$ 0.09 -0.17 $\pm$ 0.07 0.02 $\pm$ 0.09 -0.07 $\pm$ 0.08 0.13 $\pm$ 0.10 -0.10 $\pm$ 0.08 0.28 $\pm$ 0.10 -0.08 $\pm$ 0.08
IMPRESS 0.12 $\pm$ 0.10 -0.55 $\pm$ 0.07 0.11 $\pm$ 0.10 -0.54 $\pm$ 0.09 0.08 $\pm$ 0.10 -0.56 $\pm$ 0.04 0.11 $\pm$ 0.10 -0.60 $\pm$ 0.09 0.06 $\pm$ 0.10 -0.35 $\pm$ 0.09 0.11 $\pm$ 0.10 -0.10 $\pm$ 0.08 0.28 $\pm$ 0.10 -0.45 $\pm$ 0.07
Ours 0.27 $\pm$ 0.08 0.45 $\pm$ 0.06 0.27 $\pm$ 0.08 0.34 $\pm$ 0.04 0.28 $\pm$ 0.09 0.35 $\pm$ 0.06 0.29 $\pm$ 0.07 0.37 $\pm$ 0.08 0.16 $\pm$ 0.08 0.49 $\pm$ 0.06 0.19 $\pm$ 0.09 0.47 $\pm$ 0.06 0.23 $\pm$ 0.08 0.48 $\pm$ 0.07

### v4

@@ 4.1 Adversarial Perturbation Causes Latent-space Image-Prompt Mismatch
Figure 2: (a) The original causal graph representing the variable relationships in personalized diffusion model learning. Red arrows indicate false correlations introduced by adversarial perturbation. (b) Intervened causal graph with our proposed CDL. With concept extraction, we examine that CDL alone helps the model learn the right correlations for linking identifier $\mathcal{V}^{*}$ and personalized concept.
@@
$\mathcal{L}_{db}^{adv}(\mathbf{x}_{0},\bm{c}^{\mathcal{V}^{*}},\bar{\bm{x}}_{0},\bm{c};\bm{\theta})=\mathcal{L}_{\mathrm{denoise}}\left(\mathbf{x}_{0}+\bm{\delta},\bm{c}^{\mathcal{V}^{*}}\right)+\lambda\mathcal{L}_{\mathrm{denoise}}\left(\bar{\bm{x}}_{0},\bm{c}\right).$ (5)
Based on the adversarial loss in Eq. 5, with annotation of $\bm{c}^{\mathcal{V}^{*}}=\bm{c}\oplus\mathcal{V}^{*}$ where $\mathcal{V}^{*}$ denotes the embedding of the unique identifier, we build the underlying causal graph (Pearl, 2009) in (a) of Fig. 2 (See App. C.1 on the construction details) to represent the learning process of the personalized diffusion model for linking personalized identifier to instance concept. From this graph, we found that there are two false correlations (colored in red) derived from the prior prompt condition $\bar{\bm{c}}$ and instance condition $\mathcal{V}^{*}$ to the noise $\bm{\delta}$. In an ideal scenario, the adversarial perturbation $\bm{\delta}$ represents a completely relevant concept and should be independent of both the prior prompt $\bar{\bm{c}}$ and the unique identifier $\mathcal{V}^{*}$. However, during training, the model observes pairs of perturbed images $\mathbf{x}_{0}+\bm{\delta}$ and instance prompts $\mathbf{c}^{\mathcal{V}^{*}}$, leading to unintended associations between $\bm{\delta}$ and $\mathcal{V}^{*}$. To validate this, we prompt the model trained on perturbed data to generate images on two different prompts, “a photo of $\mathcal{V}^{*}$ and “a photo of $\mathcal{V}^{*}$ Person”, and visualize the results in Fig. 2. As we can see, the model erroneously attributes the noise patterns introduced by $\bm{\delta}$ to $\mathcal{V}^{*}$ and consequently generates noisy portraits for concept “$\mathcal{V}^{*}$ Person”.
To establish and reinforce the spurious path, we found that one important property that effective perturbation methods should have is the ability to cause latent-space image-prompt mismatch. That is, the images and their corresponding prompts are not semantically aligned in the latent space after the perturbation. Then thus, when learning on such pairs, it will create contradiction and force the models to dump that chaotic perturbation pattern into the rarely-appeared identifier token $\mathcal{V}^{*}$ instead of learning the clean identity behind $\bm{x}_{0}$. We infer it based on two empirical observations: i) random perturbation with the same strength does not affect the learning performance of the personalized diffusion model; ii) the generated portraits using the perturbed diffusion model usually have lower quality and larger image distortion than the slightly perturbed input images. The first observation justifies that if the perturbation does not cause a significant latent shift, then the learning of the personalized diffusion model will not be affected, while the second observation suggests that the perturbed model learns more abstract noise concepts instead of just the noise pattern in the input pixel space. We further validate this through the following experiments of latent-mismatch visualization and concept interpretation.
Figure 3: Latent 2D visualization and concept classification of images using CLIP encoders.
Specifically, using the paired CLIP encoders, we first embed the latent of clean and perturbed images and also embed the textual concept of a person with a list of prompts describing the person concept, such as “a photo of person’s face”. Then, we leverage three distinct 2D visualization techniques, including TSNE (Maaten & Hinton, 2008), Trunctued-SVD (Halko et al., 2011), and UMAP (McInnes et al., 2018) on image-prompt embedding pairs. The results in Fig. 3 suggest that adversarial perturbation indeed significantly shifts the portrait latent from its original region of the “person” concept. Moreover, we precisely split the latent space into two regions with a zero-shot CLIP-based classifier. We find that the perturbed images have a higher probability of being classified into the “noise” region instead of the “person” region in latent space. Please refer to the App. B.2 for more interpretation experiments.
These findings indicate adversarial perturbation indeed leads to latent mismatch, which is also aligned with the previous studies that reveal the fundamental vulnerability of underlying latent encoders like VAE and CLIP encoders (Akrami et al., 2020; Schlarmann et al., 2024; Xue et al., 2023). This latent mismatch creates an opportunity for shortcut learning (Geirhos et al., 2020), where the model may exploit the easier-to-learn correlation between the noise patterns and the unique identifier $\mathcal{V}^{*}$, rather than learning the true underlying concept of the person’s identity. Such shortcut learning can lead to poor generalization and vulnerability to adversarial attacks, as the model relies on spurious correlations rather than robust, semantically meaningful features.
4.2 Training Clean PDMs on Perturbed Data with Systematic Defending
To address this shortcut learning issue, we propose a systematic defense framework inspired by causal intervention (Geirhos et al., 2020), which is a widely used technique to mitigate shortcut learning in traditional machine learning tasks. Causal intervention (Kaddour et al., 2022) usually involves data argumentation or modifying the training process to disrupt spurious correlations. To mitigate shortcut learning in PDMs, we propose two key strategies: (i) Removing Noise Variables through image restoration techniques to eliminate adversarial noise and realign images with their true semantic representations; and (ii) Weakening Spurious Paths and Strengthening Causal Paths via Contrastive Decoupling Learning, which disentangles personalized concepts from noise patterns by incorporating noise tokens into prompts and leveraging clean prior data. We detail these approaches below and summarize our framework in Algorithm 1. Please refer to the App. C.2 for more details.
@@ 4.2 Training Clean PDMs on Perturbed Data with Systematic Defending
Image Purification via Image Restoration. An intuitive and effective approach to removing the direct influence of adversarial noise is to purify the input images using image restoration techniques. We view the perturbed images as degraded images in the image restoration domain (Wang et al., 2021) and leverage off-the-shelf image restoration models to convert low-quality, noisy images into high-quality, purified ones. Specifically, we use a face-oriented model named CodeFormer (Liu et al., 2023), which is trained on facial data to restore images based on latent code discretization. To further enhance the purification of non-face regions, we employ an additional diffusion-based super-resolution (SR) model. Compared to previous state-of-the-art optimization-based purification methods (Cao et al., 2024) and diffusion-based purification methods (Zhao et al., 2024), this simple yet effective pipeline yields faithful purified images with better efficiency since it only requires a single inference pass. We term this module CodeSR as it combines CodeFormer and SR in sequence.
Contrastive Decoupling Learning (CDL). To further mitigate shortcut learning, we introduce Contrastive Decoupling Learning, which aims to disentangle the learning of desired personalized concepts from undesired noise patterns. We achieve this by augmenting the prompts with additional tokens related to the noise pattern, denoted as $\mathcal{V}^{*}_{N}$, such as “XX noisy pattern”. Ideally, these newly added tokens absorb all the noise components in the image, leaving the clean, personalized concept associated with the personalized identifier $\mathcal{V}^{*}$. During training, we insert $\mathcal{V}^{*}_{N}$ into the prompt of instance data with the suffix “with XX noisy pattern”, and include the “inverse” of $\mathcal{V}^{*}_{N}$ in the prompt of class-prior data with the suffix “without XX noisy pattern”. This contrastive prompt design encourages the model to distinguish between the instance concept and noise patterns, thus weakening spurious correlations. During inference, we add the suffix “without XX noisy pattern” to the prompt input to guide the model in disregarding the learned patterns associated with $\mathcal{V}^{*}_{N}$, thereby generating images that focus on the personalized concept. Furthermore, by using classifier-free guidance (Ho & Salimans, 2022) with a negative prompt $c_{\text{neg}}=$“noisy, abstract, pattern, low quality”, we can further guide the trained model to generate high-quality images related to the learned concept. Specifically, given timestamp $t$, we perform sampling using the linear combination of the good-quality and bad-quality conditional noise estimates with guidance weight $w^{\text{neg}}=7.5$:
@@
$\tilde{\bm{\epsilon}}_{\theta}\left(\mathbf{z}_{t},\mathbf{c}\right)=(1+w^{\text{neg}})\bm{\epsilon}_{\theta}\left(\mathbf{z}_{t},\mathbf{c}^{[\mathcal{V}^{*},\bar{\mathcal{V}}_{N}^{*}]}\right)-w^{\text{neg}}\bm{\epsilon}_{\theta}\left(\mathbf{z}_{t},\tau_{\theta}(c_{\text{neg}})\right)$ (6)
Algorithm 1 Training Clean Personalized LDMs on Perturbed Data with Systematic Defense
1:Corrupted training set $X^{\prime}_{0}$, pre-trained LDM $\theta_{0}$, CodeFormer $\phi=\{\mathcal{E}_{\phi},\mathcal{D}_{\phi},\mathcal{T}_{\phi},\mathcal{C}\}$, SR model $\psi$, prior data $\bar{X}_{0}$, noise token $\mathcal{V}_{N}^{*}$, personalized identifier $\mathcal{V}^{*}$, instance prompt $c^{\mathcal{V}^{*}}$, class prompt $c$, number of generations $N_{\text{gen}}$
2:Personalized diffusion model with clean-level generation performance $\theta_{T}$
3:Step 1: Input Purification with CodeFormer and Super-resolution Model
4:CodeFormer: Predict code $\tilde{Z}_{c}=\mathcal{T}_{\phi}(\mathcal{E}_{\phi}(X^{\prime}_{0}),\mathcal{C})$; obtain high-quality restoration $\tilde{X}_{0}=\mathcal{D}_{\phi}(\tilde{Z}_{c})$
5:Super-resolution: Resize $\tilde{X}_{0}$ to $128\times 128$; apply SR model $\psi$ to obtain $\tilde{X}_{0}^{\text{purified}}$ at $512\times 512$
6:Step 2: Contrastive Decoupling Learning
7:for $i=1$ to $T$ training steps do
8:     Sample instance data $x_{i}$ from $\tilde{X}_{0}^{\text{purified}}$, and class-prior data $\bar{x}_{0}$ from $\bar{X}_{0}$
9:     Craft decoupled instance prompt $c^{\mathcal{V}^{*}}_{\text{dec}}=\texttt{concat}(c^{\mathcal{V}^{*}},\mathcal{V}_{N}^{*})$ and class-prior prompt $c_{\text{dec}}=\texttt{concat}(c,\text{``without''},\mathcal{V}_{N}^{*})$
10:     Optimize the LDM $\theta_{i}$ with standard DreamBooth loss $\mathcal{L}_{\text{db}}$triangleright Following Eq. 2
11:     $\mathcal{L}_{\text{db}}(x_{i},c^{\mathcal{V}^{*}}_{\text{dec}},\bar{x}_{0},c_{\text{dec}};\theta_{i})=\mathcal{L}_{\mathrm{denoise}}\left(x_{i},c^{\mathcal{V}^{*}}_{\text{dec}}\right)+\lambda\mathcal{L}_{\mathrm{denoise}}\left(\bar{x}_{0},c_{\text{dec}}\right)$
12:     Update LDM $\theta_{i}$ with $\nabla_{\theta_{i}}\mathcal{L}_{\text{db}}$ using AdamW optimizer on UNet Denoiser and Text Encoder
13:end for
14:Inference: Perform decoupled sampling $\{X_{\text{gen}}^{j}\}_{j=1}^{N_{\text{gen}}}$ with the trained PDM $\triangleright$ Following Eq. 6
@@ 5.1 Experimental Setup
Datasets and Metrics. Our experiments are mainly performed on the VGGFace2 (Cao et al., 2018) face dataset following (Van Le et al., 2023; Liu et al., 2024). Four identities are selected from each dataset, and we randomly pick eight images from each individual and split those images into two subsets for image protection and reference. Moreover, we also visually demonstrate the purification ability of our approach on samples from an artwork painting dataset, WikiArt (Saleh & Elgammal, 2015), and the CelebA (Liu et al., 2015). For the metric, we evaluate the generated images in terms of their semantic-related quality and graphical aesthetic quality. For the semantic-related score, we compute the cosine similarity between the embedding of generated images and reference images, which we term the Identity Matching Similarity (IMS) score. We reported the weighted averaged IMS score by employing two face embedding extractors, including antelopev2 model from InsightFace library (Deng et al., 2020) following IP-adapter (Ye et al., 2023) and VGG-Net (Simonyan & Zisserman, 2014) from Deepface library (Taigman et al., 2014) following (Van Le et al., 2023). The IMS score is computed via a weighted sum: IMS$=\lambda\text{IMS}_{\text{IP}}+(1-\lambda)\text{IMS}_{\text{VGG}}$, where $\lambda$ is set as 0.7. For the graphical quality $Q$, we report the average of two metrics: i) LIQE (Zhang et al., 2023) (with re-normalization to $[-1,+1]$); ii) CLIP-IQAC following (Liu et al., 2024), which is based on CLIP-IQA (Wang et al., 2023) with class label. See App. A.1 for details.
@@
Purification Baselines and Perturbation Methods. For purification baselines, we consider both model-free and diffusion-based approaches. The model-free methods include ❶ Gaussian Filtering, which reduces noise and detail using a Gaussian kernel; ❷ Total Variation Minimization (TVM), which reconstructs images by minimizing the difference between original and reconstructed images while enforcing smoothness; and ❸ JPEG Compression, which reduces image file size by transforming images into a compressed format. The diffusion-based methods include ❹ (Pixel)DiffPure (Nie et al., 2022), which leverages pretrained pixel-space diffusion models to smooth adversarial noise with small-step SDEdit process (Meng et al., 2021); ❺ LatentDiffPure, which is developed in the paper similar as DiffPure but with LDM as a purifier (two variants w/ and w/o prompting); ❻ DDSPure (Carlini et al., 2022), which finds an optimal timestamp for adversarial purification with SDEdit process; ❼ GrIDPure (Zheng et al., 2023), which further conducts iterative DiffPure with small steps with grid-based splitting to improve structure similarity; and ❽ IMPRESS (Cao et al., 2024), which purifies by optimizing latent consistency with visual similarity constraints. For protective perturbation, we consider six of existing SoTA approaches, including perturbation crafted with bi-level optimization, such as FSMG, ASPL, EASPL (Van Le et al., 2023), MetaCloak (Liu et al., 2024), and perturbations crafted with adversarial perturbation with fixed models, such as AdvDM (Liang et al., 2023), PhotoGuard (Salman et al., 2023), and Glaze (Shan et al., 2023). For each setting, we set the perturbation to be ASPL by default. We set the $\ell_{\infty}$ radius to 11/255 with a six-step PGD step size of 1/255 by default following (Van Le et al., 2023). See App. A.2 for more details.
@@
[…]
Clean -0.13 0.15 -0.13 0.15 -0.13 0.15 -0.13 0.15 -0.13 0.15 -0.13 0.15 -0.13 0.15
Perturbed -0.43 -0.54 -0.67 -0.52 -0.62 -0.50 -0.35 -0.53 -0.27 -0.36 -0.18 -0.24 -0.28 -0.28
Gaussian F. -0.19 -0.55 -0.20 -0.57 -0.17 -0.58 -0.07 -0.63 -0.11 -0.57 -0.23 -0.53 -0.18 -0.54
JPEG -0.15 -0.41 -0.21 -0.52 -0.27 -0.50 -0.34 -0.38 -0.15 -0.02 -0.13 0.07 -0.19 -0.03
TVM -0.12 -0.65 -0.16 -0.66 -0.10 -0.67 -0.11 -0.69 -0.12 -0.65 -0.15 -0.64 -0.11 -0.66
PixelDiffPure -0.41 -0.57 -0.43 -0.54 -0.57 -0.61 -0.28 -0.58 -0.40 -0.55 -0.25 -0.55 -0.41 -0.59
L.DiffPure-$\emptyset$ -0.07 -0.47 -0.36 -0.59 -0.22 -0.49 -0.52 -0.43 -0.55 -0.24 -0.12 -0.40 -0.38 -0.42
L.DiffPure -0.25 -0.45 -0.31 -0.61 -0.30 -0.46 -0.31 -0.51 -0.57 -0.30 -0.25 -0.47 -0.41 -0.47
DDSPure -0.15 -0.34 -0.05 -0.38 -0.08 -0.39 -0.16 -0.49 -0.19 -0.43 -0.12 -0.37 -0.22 -0.41
GrIDPure -0.10 -0.20 -0.21 -0.16 -0.13 -0.25 -0.23 -0.25 -0.09 -0.18 -0.03 -0.22 -0.24 -0.13
IMPRESS -0.15 -0.53 -0.16 -0.49 -0.29 -0.64 -0.34 -0.29 -0.34 -0.34 -0.16 -0.21 -0.10 -0.43
Ours 0.23 0.65 0.09 0.62 0.09 0.63 0.38 0.58 0.29 0.67 0.24 0.63 0.31 0.66